# Heterogeneous Complementary Distillation

### v1

@@ Introduction
Knowledge distillation (KD), proposed by Hinton et al. (hinton2015distilling), aligns the soft label distributions of a complex teacher model (abbreviated teacher) and a compact student model (abbreviated student) using Kullback-Leibler (KL) divergence (kullback1951information). This technique enables the student’s performance to approach that of the teacher’s. Due to its simplicity and its ability to preserve the original architecture, KD is widely applied to resource-constrained computer vision tasks, such as image classification (he2016deep; yang2022cross; zheng2025hierarchical), object detection (li2024detkds; zhao2024detrs; li2023kd) and semantic segmentation (zou2024segment; liang2023open). However, traditional variants of KD, such as FitNets (romero2014fitnets), DKD (zhao2022decoupled), and RKD (park2019relational), mainly focus on distillation between homogeneous architectures (e.g., ResNet34 to ResNet18). This limits the improvement of the student’s performance due to the lack of knowledge diversity in the KD strategy. In contrast, Vision Transformers (ViTs) (dosovitskiy2020image; tolstikhin2021mlp), with stronger representational capacity, provide new avenues for heterogeneous architecture (or cross-architecture) distillation (liu2022cross) and allow the student to acquire diverse feature representations from the teacher. Intuitively, heterogeneous architecture distillation, such as ViT-to-ResNet18, leverages the complementary information of inductive biases from the teacher and the student. As a result, it enhances the student’s robustness and generalization beyond the limitations of homogeneous distillation methods. However, when traditional feature-based distillation methods (romero2014fitnets; park2019relational) in homogeneous settings are applied to heterogeneous settings, distillation result often fails due to mismatched spatial feature representations (e.g., differences in receptive field size), which results in the degradation of the student’s performance.
@@
To overcome the heterogeneous feature representation gap, existing methods, such as OFA-KD (hao2023one), adopt a one-for-all strategy at each student stage to map intermediate representations to the logits space and align them with the teacher’s logits. However, relying solely on logits alignment between teacher and student fails to leverage their complementary strengths, thus limiting potential performance gains. Similarly, PAT proposes a region-aware attention (RAA) module (lin2025feature) that restructures student intermediate features to align with teacher representations, improving student’s performance but incurring high computational costs and complex designs. These limitations highlight the challenge of transferring heterogeneous knowledge effectively, primarily due to the complex spatial feature patterns that students struggle to fully capture in heterogeneous distillation. In particular, over-reliance on the teacher’s logits, as seen in OFA-KD, may cause the student to overlook its own feature extraction strengths. Therefore, our research focuses on tailoring logit knowledge to integrate the feature representations from both teacher’s and the student’s architectural strengths, facilitating more effective knowledge transfer in heterogeneous distillation.
@@
Homogeneous Architecture Knowledge Distillation. Knowledge distillation (KD) transfers the teacher’s “dark knowledge” to a compact student, enabling the student to achieve performance comparable to the teacher’s while reducing the model size. KD methods are broadly categorized into feature-based KD (li2023curriculum; yang2022focal; zhao2024detrs; romero2014fitnets; liu2023norm; tian2019contrastive; li2023automated; chen2021distilling; liu2023functionconsistent; dong2023diswot; guo2023class; huang2024knowledge; kim2024do; yang2022masked; yang2021knowledge; wang2019distilling; li2022shadow) and logit-based KD (zhao2022decoupled; yang2023knowledge; li2022asymmetric; huang2022knowledge; xu2024improving; luo2024scale; Auxiliary; zheng2024knowledge; Sun2024Logit; jin2023multi). Feature-based KD methods use the teacher’s intermediate features to improve the student’s feature representations. For example, FitNets (romero2014fitnets) aligns intermediate features across stages using mean squared error. NORM (liu2023norm) enables efficient many-to-one feature matching via expanded student representations. CRD (tian2019contrastive) incorporates contrastive learning to improve knowledge transfer efficiency. ReviewKD (chen2021distilling) introduces a review learning mechanism that improves the efficiency of knowledge transfer from the teacher. Despite their effectiveness, these methods often incur high computational and memory cost, prompting a shift toward logit-based distillation approaches. For instance, DKD (zhao2022decoupled) decouples logit knowledge into target and non-target class KD. LDRLD (xu2025local) mines fine-grained logit relationships to improve the student’s performance. IKD (wang2025debiased) uses intra-class distillation to mitigate prediction biases within the same class. WTTM (zheng2024knowledge) refines the temperature scaling in teacher matching. RLD (sun2024knowledge) uses labeling information to dynamically refine teacher knowledge. TeKAP (hossain2025single) employs multi-perspective teacher knowledge via random sampling. Following the success of KD methods in homogeneous architecture, researchers have shifted to address architectural disparities between the teacher and the student.
@@
Heterogeneous Architecture Knowledge Distillation. Architectural disparities between the teacher and student often limit the effectiveness of homogeneous distillation methods in heterogeneous settings. To tackle this challenge, researchers have proposed various methods (hao2023one; li2024tas; lin2025feature; zhao2023cumulative; zhou2025all; huang2025distilling; lee2025customkd; zhang2025cross). For example, MLDR-KD (yang2025multilevel) proposes a novel heterogeneous relational KD framework that retains dark knowledge while boosting confidence in the correct target. RSD (zhang2025cross) extracts architecture-agnostic knowledge by reducing redundant architecture-specific information. PAT (lin2025feature) introduces a region-aware attention (RAA) module that aligns teacher representations by restructuring multi-layer student features. LFCC (wu2024aligning) uses a multi-scale low-pass filter and its learnable derivative to compress low-frequency components of teacher and student features, reducing representational disparity and mitigating the impact of spatial noise. TCS (zhou2025all) tailors a coordinate system to transfer “dark knowledge” from a task-agnostic teacher to task-specific student networks. HeteroAKD (huang2025distilling) transforms heterogeneous knowledge into the logit space, mitigating architecture-specific influences. CustomKD (lee2025customkd) customizes the generalized features of large vision foundation models (LVFMs) for the student, minimizing spatial representation differences. However, methods like HeteroAKD, LFCC, and PAT, which rely on all intermediate stages, incur high computational costs and decrease the efficiency of knowledge transfer. Consequently, heterogeneous KD approaches require further optimization to overcome these limitations. To address these challenges, our proposed method integrates complementary feature mapping from the teacher and student, sub-logit decoupled distillation, and orthogonal constraints to enhance efficiency and effectiveness.
@@
KD was originally proposed by Hinton et al. (hinton2015distilling). By minimizing the output probability distributions of the teacher and the student via KL divergence, KD transfers the complex teacher’s “dark knowledge” to the compact student. Specifically, the KD loss, $\mathcal{L}_{KL}$, uses KL divergence to measure the difference between the soft label distributions of the student and the teacher, while the cross-entropy loss, $\mathcal{L}_{CE}$, supervises the student’s classification performance. The total distillation loss combines two terms as follows:
@@
Heterogeneous architectures, such as convolutional neural networks (CNNs), vision transformers (ViTs), and MLP-Mixers, generate distinct spatial feature representations due to their inherent inductive biases. For instance, CNNs emphasize locality and translation invariance, whereas ViTs excel at modeling global contextual relationships. Direct alignment of the intermediate feature representations using FitNets (romero2014fitnets) between a student (e.g., CNN) and a teacher (e.g., ViT) often results in failure due to distinct representations. This consequently hinders effective knowledge transfer and degrades student performance. Intuitively, heterogeneous distillation can enrich knowledge diversity by leveraging differences in feature representation between architectures, thereby improving the student’s performance. However, traditional homogeneous distillation methods, such as feature alignment in FitNets or logit distillation in KD, fail to address these representation disparities, leading to suboptimal student performance. As a result, they cannot fully exploit the inductive bias information of each architecture to enhance the student’s performance. While existing heterogeneous distillation methods attempt to address this issue, they still have limitations. For example, OFA’s single-stage logit mapping oversimplifies the process, limiting the student’s feature representation capacity. Conversely, PAT’s RAA mitigates feature representation misalignment effectively but incurs significant computational overhead. To address these issues, The proposed Heterogeneous Complementary Distillation (HCD) method. The HCD integrates the teacher’s deep feature representations with the student’s intermediate feature representations and enhances the student’s performance.
@@
Datasets. We evaluate the performance of the HCD method using the CIFAR-100 (krizhevsky2009learning) and ImageNet-1K (deng2009imagenet) datasets. Specifically, CIFAR-100 consists of 100 classes, with a total of 50,000 samples, including 500 training images and 100 test images per class, each with a resolution of $\times32$ pixels. To ensure consistency with our model’s input requirements in heterogeneous distillation, we resize all CIFAR-100 images to $\times224$ pixels before training. In contrast, ImageNet-1K contains 1,000 classes, with approximately 1.28 million training images and 50,000 validation images, each with a resolution of $\times224$ pixels. CUB-200-2021 (wah2011caltech) consists of 11,788 images from 200 bird species, widely used for fine-grained classification with annotations like bounding boxes and keypoints. FGVC-Aircraft (maji2013fine) consists of 10,000 images across 100 aircraft models, used for fine-grained classification tasks.
@@
Image Classification Results on Fine-grained datasets. To further evaluate the performance of HCD and assess its generalizability, we also extended the experiments to fine-grained datasets. Leveraging recent advancements in the powerful feature representation capabilities of vision foundation models (VFMs), we employed the visual encoders of DINOv2 (oquab2024dinov) and SigLIP2 (tschannen2025siglip) as teacher models, with various CNNs acting as student models. As shown in Tables 4 and 5, compared to exsiting KD methods (e.g., DKD,LDRLD), HCD achieves superior performance on the FGVC-Aircraft. This results demonstrate that HCD effectively uses heterogeneous complementary information to enhance feature representation in transfer learning for VFMs. However, the FGVC-Aircraft images exhibit distinct category differences, while CUB-200 presents greater learning difficulty due to its fine-grained features, resulting in better performance in heterogeneous distillation for the former in Tables 4 and 5. However, HCD outperforms LDRLD slightly on the CUB200 dataset, likely due to the fine-grained nature of the task, where LDRLD excels in capturing inter-class relationships.
@@ Acknowledgments
Baselines. To show the performance of our proposed HCD framework, we benchmark it against leading knowledge distillation techniques. In particular, for feature-based distillation, we choose FitNets (romero2014fitnets), AT (komodakis2017paying), ReviewKD (chen2021distilling), OFD (heo2019comprehensive), FCFD (liu2023functionconsistent), CAT-KD (guo2023class), RKD (park2019relational), CC (peng2019correlation), CRD (tian2019contrastive), RSD (zhang2025cross), and PAT (lin2025feature) . For logit-based distillation, we choose KD (hinton2015distilling), DKD (zhao2022decoupled), IPWD (niu2022respecting), CTKD (li2023curriculum), SDD (luo2024scale), LSKD (Sun2024Logit), DIST (huang2022knowledge), OFA-KD (hao2023one), WTTM (zheng2024knowledge), RKKD (guan2025enhancing), LDRLD (xu2025local), RLD (sun2024knowledge), and TeKAP (hossain2025single).
@@
Models. To evaluate the effectiveness of our HCD framework in heterogeneous knowledge distillation, we employ various teacher-student architectures, including Convolutional Neural Networks (CNNs), Vision Transformers (ViTs) and MLP-Mixers. Specifically, we evaluate the following architectures on the CIFAR-100 and ImageNet-1K: ResNet-50 (he2016deep),ViT (dosovitskiy2020image), Swin-Tiny (liu2021swin), MLP-Mixer-B/16 (tolstikhin2021mlp), DeiT-Tiny (touvron2021training), ResNet-18 (he2016deep), and MobileNetV1 (howard2017mobilenets), with larger models (e.g., ResNet-50, Swin-Tiny) as teachers and smaller models (e.g., ResNet-18, MobileNetV1) as students. For training fine-grained datasets, we use the encoder of the visual base model as the teacher model, specifically ViT-L and ViT-S from DINOv2 (oquab2024dinov), ViT-L and ViT-B from SigLIP2 (tschannen2025siglip). We use some CNNs as student model, such as MobileNetV2 (sandler2018mobilenetv2), VGG (simonyan2014very), ResNet20, ShuffleNetV2 (ma2018shufflenet).
@@
Training details: (1) We employ different optimizers for various student model architectures. Specifically, we train all CNN student models (e.g., ResNet-18, MobileNetV2) using SGD, while for Vision Transformers (ViTs) (e.g., Swin-Tiny) and MLP-Mixers (e.g., MLP-Mixer-B/16), we use the AdamW optimizer. We train all models for 300 epochs on the CIFAR-100 dataset and we train CNN students for 100 epochs on the ImageNet-1K dataset. For more detailed parameter configuration, see OFA-KD (hao2023one). (2) The batch size of 256 is used for same-architecture on the ImageNet-1K dataset, and the initial learning rate is set to 0.1. A linear warm-up is applied during the first 10 epochs. The learning rate is then reduced by a factor of 10 at the 30th, 60th, and 90th epochs. For more detailed parameter configuration, see DKD (zhao2022decoupled). For the student model ResNet18 on the CIFAR100 dataset, the settings are $\lambda$=1.0, $\omega$=10.0, $\beta$=8.0, and temperature $\tau$=4.0. For the student model MobileNetV2 on the CIFAR100 dataset, the settings are $\lambda$=1.0, $\omega$=1.0, $\beta$=2.0, and temperature $\tau$=4.0. For the parameter settings of heterogeneous distillation on ImageNet-1K, see Table 8 and same architecture distillation on ImageNet-1K, see Table 9. (3) We fine-tune the linear classification heads of DINOv2 and SigLIP2 on the CUB-200-2011 and FGVC-Aircraft datasets for 20 epochs, with a batch size of 128 and a learning rate of 0.1, serving as the teacher models (ViT-S,ViT-L,ViT-B). The detailed parameter settings for training the student models are similar to those used in CRD (tian2019contrastive) on the CIFAR100 dataset.

### v2

@@ Introduction
Knowledge distillation (KD), proposed by Hinton et al. (Hinton et al. 2014), aligns the soft label distributions of a complex teacher model (abbreviated teacher) and a compact student model (abbreviated student) using Kullback-Leibler (KL) divergence (Kullback and Leibler 1951). This technique enables the student’s performance to approach that of the teacher’s. Due to its simplicity and its ability to preserve the original architecture, KD is widely applied to resource-constrained computer vision tasks, such as image classification (He et al. 2016; Yang et al. 2022a; Zheng et al. 2025), object detection (Li et al. 2024b; Zhao et al. 2024; Li et al. 2023a) and semantic segmentation (Zou et al. 2024; Liang et al. 2023). However, traditional variants of KD, such as FitNets (Romero et al. 2015), DKD (Zhao et al. 2022), and RKD (Park et al. 2019), mainly focus on distillation between homogeneous architectures (e.g., ResNet34 to ResNet18). This limits the improvement of the student’s performance due to the lack of knowledge diversity in the KD strategy. In contrast, Vision Transformers (ViTs) (Dosovitskiy et al. 2021; Tolstikhin et al. 2021), with stronger representational capacity, provide new avenues for heterogeneous architecture (or cross-architecture) distillation (Liu et al. 2022) and allow the student to acquire diverse feature representations from the teacher. Intuitively, heterogeneous architecture distillation, such as ViT-to-ResNet18, leverages the complementary information of inductive biases from the teacher and the student. As a result, it enhances the student’s robustness and generalization beyond the limitations of homogeneous distillation methods. However, when traditional feature-based distillation methods (Romero et al. 2015; Park et al. 2019) in homogeneous settings are applied to heterogeneous settings, distillation result often fails due to mismatched spatial feature representations (e.g., differences in receptive field size), which results in the degradation of the student’s performance.
@@
To overcome the heterogeneous feature representation gap, existing methods, such as OFA-KD (Hao et al. 2023), adopt a one-for-all strategy at each student stage to map intermediate representations to the logits space and align them with the teacher’s logits. However, relying solely on logits alignment between teacher and student fails to leverage their complementary strengths, thus limiting potential performance gains. Similarly, PAT proposes a region-aware attention (RAA) module (Lin et al. 2025) that restructures student intermediate features to align with teacher representations, improving student’s performance but incurring high computational costs and complex designs. These limitations highlight the challenge of transferring heterogeneous knowledge effectively, primarily due to the complex spatial feature patterns that students struggle to fully capture in heterogeneous distillation. In particular, over-reliance on the teacher’s logits, as seen in OFA-KD, may cause the student to overlook its own feature extraction strengths. Therefore, our research focuses on tailoring logit knowledge to integrate the feature representations from both teacher’s and the student’s architectural strengths, facilitating more effective knowledge transfer in heterogeneous distillation.
@@
Homogeneous Architecture Knowledge Distillation. Knowledge distillation (KD) transfers the teacher’s “dark knowledge” to a compact student, enabling the student to achieve performance comparable to the teacher’s while reducing the model size. KD methods are broadly categorized into feature-based KD (Li et al. 2023c; Yang et al. 2022b; Zhao et al. 2024; Romero et al. 2015; Liu et al. 2023b; Tian et al. 2020; Li et al. 2023b; Chen et al. 2021; Liu et al. 2023a; Dong et al. 2023; Guo et al. 2023; Huang et al. 2023; Kim et al. 2024; Yang et al. 2022c; yang2021knowledge; Wang et al. 2019; Li and Jin 2022) and logit-based KD (Zhao et al. 2022; Yang et al. 2023; Li et al. 2022; Huang et al. 2022; Xu et al. 2024; Wei et al. 2024; Peng et al. 2024; Zheng and Yang 2024; Sun et al. 2024; Jin et al. 2023). Feature-based KD methods use the teacher’s intermediate features to improve the student’s feature representations. For example, FitNets (Romero et al. 2015) aligns intermediate features across stages using mean squared error. NORM (Liu et al. 2023b) enables efficient many-to-one feature matching via expanded student representations. CRD (Tian et al. 2020) incorporates contrastive learning to improve knowledge transfer efficiency. ReviewKD (Chen et al. 2021) introduces a review learning mechanism that improves the efficiency of knowledge transfer from the teacher. Despite their effectiveness, these methods often incur high computational and memory cost, prompting a shift toward logit-based distillation approaches. For instance, DKD (Zhao et al. 2022) decouples logit knowledge into target and non-target class KD. LDRLD (Xu et al. 2025) mines fine-grained logit relationships to improve the student’s performance. IKD (Wang et al. 2025) uses intra-class distillation to mitigate prediction biases within the same class. WTTM (Zheng and Yang 2024) refines the temperature scaling in teacher matching. RLD (Sun et al. 2025) uses labeling information to dynamically refine teacher knowledge. TeKAP (Hossain et al. 2025) employs multi-perspective teacher knowledge via random sampling. Following the success of KD methods in homogeneous architecture, researchers have shifted to address architectural disparities between the teacher and the student.
@@
Heterogeneous Architecture Knowledge Distillation. Architectural disparities between the teacher and student often limit the effectiveness of homogeneous distillation methods in heterogeneous settings. To tackle this challenge, researchers have proposed various methods (Hao et al. 2023; Li et al. 2024a; Lin et al. 2025; Zhao et al. 2023; Zhou et al. 2025; Huang et al. 2025; Lee et al. 2025; Zhang et al. 2025). For example, MLDR-KD (Yang et al. 2025) proposes a novel heterogeneous relational KD framework that retains dark knowledge while boosting confidence in the correct target. RSD (Zhang et al. 2025) extracts architecture-agnostic knowledge by reducing redundant architecture-specific information. PAT (Lin et al. 2025) introduces a region-aware attention (RAA) module that aligns teacher representations by restructuring multi-layer student features. LFCC (Wu et al. 2024) uses a multi-scale low-pass filter and its learnable derivative to compress low-frequency components of teacher and student features, reducing representational disparity and mitigating the impact of spatial noise. TCS (Zhou et al. 2025) tailors a coordinate system to transfer “dark knowledge” from a task-agnostic teacher to task-specific student networks. HeteroAKD (Huang et al. 2025) transforms heterogeneous knowledge into the logit space, mitigating architecture-specific influences. CustomKD (Lee et al. 2025) customizes the generalized features of large vision foundation models (LVFMs) for the student, minimizing spatial representation differences. However, methods like HeteroAKD, LFCC, and PAT, which rely on all intermediate stages, incur high computational costs and decrease the efficiency of knowledge transfer. Consequently, heterogeneous KD approaches require further optimization to overcome these limitations. To address these challenges, our proposed method integrates complementary feature mapping from the teacher and student, sub-logit decoupled distillation, and orthogonal constraints to enhance efficiency and effectiveness.
@@
KD was originally proposed by Hinton et al. (Hinton et al. 2014). By minimizing the output probability distributions of the teacher and the student via KL divergence, KD transfers the complex teacher’s “dark knowledge” to the compact student. Specifically, the KD loss, $\mathcal{L}_{KL}$, uses KL divergence to measure the difference between the soft label distributions of the student and the teacher, while the cross-entropy loss, $\mathcal{L}_{CE}$, supervises the student’s classification performance. The total distillation loss combines two terms as follows:
@@
Heterogeneous architectures, such as convolutional neural networks (CNNs), vision transformers (ViTs), and MLP-Mixers, generate distinct spatial feature representations due to their inherent inductive biases. For instance, CNNs emphasize locality and translation invariance, whereas ViTs excel at modeling global contextual relationships. Direct alignment of the intermediate feature representations using FitNets (Romero et al. 2015) between a student (e.g., CNN) and a teacher (e.g., ViT) often results in failure due to distinct representations. This consequently hinders effective knowledge transfer and degrades student performance. Intuitively, heterogeneous distillation can enrich knowledge diversity by leveraging differences in feature representation between architectures, thereby improving the student’s performance. However, traditional homogeneous distillation methods, such as feature alignment in FitNets or logit distillation in KD, fail to address these representation disparities, leading to suboptimal student performance. As a result, they cannot fully exploit the inductive bias information of each architecture to enhance the student’s performance. While existing heterogeneous distillation methods attempt to address this issue, they still have limitations. For example, OFA’s single-stage logit mapping oversimplifies the process, limiting the student’s feature representation capacity. Conversely, PAT’s RAA mitigates feature representation misalignment effectively but incurs significant computational overhead. To address these issues, The proposed Heterogeneous Complementary Distillation (HCD) method. The HCD integrates the teacher’s deep feature representations with the student’s intermediate feature representations and enhances the student’s performance.
@@
Datasets. We evaluate the performance of the HCD method using the CIFAR-100 (Krizhevsky et al. 2009) and ImageNet-1K (Deng et al. 2009) datasets. Specifically, CIFAR-100 consists of 100 classes, with a total of 50,000 samples, including 500 training images and 100 test images per class, each with a resolution of $\times32$ pixels. To ensure consistency with our model’s input requirements in heterogeneous distillation, we resize all CIFAR-100 images to $\times224$ pixels before training. In contrast, ImageNet-1K contains 1,000 classes, with approximately 1.28 million training images and 50,000 validation images, each with a resolution of $\times224$ pixels. CUB-200-2021 (Wah et al. 2011) consists of 11,788 images from 200 bird species, widely used for fine-grained classification with annotations like bounding boxes and keypoints. FGVC-Aircraft (Maji et al. 2013) consists of 10,000 images across 100 aircraft models, used for fine-grained classification tasks.
@@
Image Classification Results on Fine-grained datasets. To further evaluate the performance of HCD and assess its generalizability, we also extended the experiments to fine-grained datasets. Leveraging recent advancements in the powerful feature representation capabilities of vision foundation models (VFMs), we employed the visual encoders of DINOv2 (Oquab et al. 2024) and SigLIP2 (Tschannen et al. 2025) as teacher models, with various CNNs acting as student models. As shown in Tables 4 and 5, compared to exsiting KD methods (e.g., DKD,LDRLD), HCD achieves superior performance on the FGVC-Aircraft. This results demonstrate that HCD effectively uses heterogeneous complementary information to enhance feature representation in transfer learning for VFMs. However, the FGVC-Aircraft images exhibit distinct category differences, while CUB-200 presents greater learning difficulty due to its fine-grained features, resulting in better performance in heterogeneous distillation for the former in Tables 4 and 5. However, HCD outperforms LDRLD slightly on the CUB200 dataset, likely due to the fine-grained nature of the task, where LDRLD excels in capturing inter-class relationships.
@@ Acknowledgments
Baselines. To show the performance of our proposed HCD framework, we benchmark it against leading knowledge distillation techniques. In particular, for feature-based distillation, we choose FitNets (Romero et al. 2015), AT (Komodakis and Zagoruyko 2017), ReviewKD (Chen et al. 2021), OFD (Heo et al. 2019), FCFD (Liu et al. 2023a), CAT-KD (Guo et al. 2023), RKD (Park et al. 2019), CC (Peng et al. 2019), CRD (Tian et al. 2020), RSD (Zhang et al. 2025), and PAT (Lin et al. 2025) . For logit-based distillation, we choose KD (Hinton et al. 2014), DKD (Zhao et al. 2022), IPWD (Niu et al. 2022), CTKD (Li et al. 2023c), SDD (Wei et al. 2024), LSKD (Sun et al. 2024), DIST (Huang et al. 2022), OFA-KD (Hao et al. 2023), WTTM (Zheng and Yang 2024), RKKD (Guan et al. 2025), LDRLD (Xu et al. 2025), RLD (Sun et al. 2025), and TeKAP (Hossain et al. 2025).
@@
Models. To evaluate the effectiveness of our HCD framework in heterogeneous knowledge distillation, we employ various teacher-student architectures, including Convolutional Neural Networks (CNNs), Vision Transformers (ViTs) and MLP-Mixers. Specifically, we evaluate the following architectures on the CIFAR-100 and ImageNet-1K: ResNet-50 (He et al. 2016),ViT (Dosovitskiy et al. 2021), Swin-Tiny (Liu et al. 2021), MLP-Mixer-B/16 (Tolstikhin et al. 2021), DeiT-Tiny (Touvron et al. 2021), ResNet-18 (He et al. 2016), and MobileNetV1 (Howard et al. 2017), with larger models (e.g., ResNet-50, Swin-Tiny) as teachers and smaller models (e.g., ResNet-18, MobileNetV1) as students. For training fine-grained datasets, we use the encoder of the visual base model as the teacher model, specifically ViT-L and ViT-S from DINOv2 (Oquab et al. 2024), ViT-L and ViT-B from SigLIP2 (Tschannen et al. 2025). We use some CNNs as student model, such as MobileNetV2 (Sandler et al. 2018), VGG (Simonyan and Zisserman 2015), ResNet20, ShuffleNetV2 (Ma et al. 2018).
@@
Training details: (1) We employ different optimizers for various student model architectures. Specifically, we train all CNN student models (e.g., ResNet-18, MobileNetV2) using SGD, while for Vision Transformers (ViTs) (e.g., Swin-Tiny) and MLP-Mixers (e.g., MLP-Mixer-B/16), we use the AdamW optimizer. We train all models for 300 epochs on the CIFAR-100 dataset and we train CNN students for 100 epochs on the ImageNet-1K dataset. For more detailed parameter configuration, see OFA-KD (Hao et al. 2023). (2) The batch size of 256 is used for same-architecture on the ImageNet-1K dataset, and the initial learning rate is set to 0.1. A linear warm-up is applied during the first 10 epochs. The learning rate is then reduced by a factor of 10 at the 30th, 60th, and 90th epochs. For more detailed parameter configuration, see DKD (Zhao et al. 2022). For the student model ResNet18 on the CIFAR100 dataset, the settings are $\lambda$=1.0, $\omega$=10.0, $\beta$=8.0, and temperature $\tau$=4.0. For the student model MobileNetV2 on the CIFAR100 dataset, the settings are $\lambda$=1.0, $\omega$=1.0, $\beta$=2.0, and temperature $\tau$=4.0. For the parameter settings of heterogeneous distillation on ImageNet-1K, see Table 8 and same architecture distillation on ImageNet-1K, see Table 9. (3) We fine-tune the linear classification heads of DINOv2 and SigLIP2 on the CUB-200-2011 and FGVC-Aircraft datasets for 20 epochs, with a batch size of 128 and a learning rate of 0.1, serving as the teacher models (ViT-S,ViT-L,ViT-B). The detailed parameter settings for training the student models are similar to those used in CRD (Tian et al. 2020) on the CIFAR100 dataset.